# A Setwise Approach for Effective and Highly Efficient Zero-shot Ranking with Large Language Models

### v1

@@ 5.1. Effectiveness Results
In regards to ranking effectiveness, it is notable that all LLM-based zero-shot ranking approaches demonstrate a significant improvement over the initial BM25 ranking. The only exception to this trend is the pointwise.qlm approach on DL2019 across all models and DL2020 with the Flan-t5-xxl model. Interestingly, as the LLM size increases, the effectiveness of pointwise.qlm decreases. This finding is particularly unexpected, given the common assumption that larger LLMs tend to be more effectiveness. On the other hand, pointwise.yes_no method achieved a decent NDCG@10 score with Flan-t5-large when compared to other methods. However, effectiveness also did not increase as model size increased. These unexpected results for both Pointwise methods might be attributed to the requirement of a more refined model output calibration process, ensuring their suitability for comparison and sorting across different documents (Qin et al., 2023). The Listwise approaches (listwise.generation) are far less effective when tested with Flan-t5-large and Flan-t5-xl. However, listwise.generation shows some improvement with Flan-t5-xxl. These results may be attributed to the fact that generating a ranking list requires fine-grained relevance preferences across multiple documents, a task that may exceed the capabilities of smaller models. In contrast, the listwise.likelihood approach, empowered by our Setwise prompt, markedly enhances the ranking effectiveness of the Listwise approach, even when utilizing smaller models. We acknowledge however that listwise.likelihood requires access to the model output logits, whereas listwise.generation does not. In the case of Pairwise and Setwise approaches, they consistently exhibit good ranking effectiveness across various model sizes and datasets.
@@
In Table 3, we present the zero-shot ranking effectiveness of all methods (with the exception of pairwise.allpair due to its computationally intensive nature) across 9 widely-used BEIR datasets. Notably, we identify several different trends that deviate from observations made on the TREC DL datasets. Firstly, pointwise.qlm exhibits a slightly higher average NDCG@10 score compared to pointwise.yes_no. Moreover, the effectiveness of pointwise.qlm remains stable even as the model size increases. Secondly, listwise.generation demonstrates comparable effectiveness to listwise.likelihood, with the majority of gains obtained in the Touche dataset, where other methods perform worse. Lastly, both Pairwise and Setwise methods that leverage the bubble sort algorithm consistently demonstrate higher average NDCG@10 compared to when they utilize the heap sort algorithm, regardless of the model size. Overall, the variety of results we observe across different experimental settings shows the importance of not drawing conclusions about effectiveness from single datasets or model sizes.
a | BM25 | .595 | .322 | .442 | .318 | .679 | .331 | .395 | .407 | .436 | |
Flan-t5-large | b | pointwise.qlm | .664 | .322 | .260 | .305 | .644 | .314 | .413 | .439 | .420 |
c | pointwise.yes_no | .664 | .308 | .238 | .296 | .504 | .275 | .346 | .456 | .386 | |
d | listwise.generation | .692 | .333 | .441 | .391 | .650 | .343 | .428 | .441 | .465 | |
e | listwise.likelihood | .756 | .334 | .327 | .444 | .639 | .308 | .453 | .475 | .467 | |
f | pairwise.heapsort | .761 | .336 | .318 | .414 | .671 | .325 | .440 | .402 | .458 | |
g | pairwise.bubblesort | .714 | .341 | .447 | .416 | .700 | .361 | .440 | .439 | .482 | |
h | setwise.heapsort | .768 | .325 | .303 | .413 | .620 | .319 | .439 | .462 | .456 | |
i | setwise.bubblesort | .761 | .338 | .394 | .441 | .636 | .351 | .447 | .497 | .483 | |
Flan-t5-xl | b | pointwise.qlm | .679 | .330 | .216 | .310 | .696 | .299 | .422 | .427 | .422 |
c | pointwise.yes_no | .698 | .331 | .269 | .273 | .553 | .297 | .413 | .479 | .414 | |
d | listwise.generation | .650 | .334 | .451 | .366 | .694 | .349 | .437 | .475 | .470 | |
e | listwise.likelihood | .736 | .360 | .310 | .449 | .686 | .320 | .472 | .526 | .482 | |
f | pairwise.heapsort | .778 | .355 | .303 | .417 | .711 | .317 | .471 | .550 | .488 | |
g | pairwise.bubblesort | .763 | .359 | .400 | .432 | .734 | .353 | .485 | .553 | .510 | |
h | setwise.heapsort | .757 | .352 | .283 | .428 | .677 | .314 | .465 | .520 | .475 | |
i | setwise.bubblesort | .756 | .353 | .330 | .438 | .691 | .362 | .497 | .537 | .496 | |
Flan-t5-xxl | b | pointwise.qlm | .707 | .342 | .188 | .324 | .712 | .307 | .431 | .440 | .431 |
c | pointwise.yes_no | .691 | .322 | .240 | .305 | .623 | .274 | .392 | .515 | .420 | |
d | listwise.generation | .664 | .344 | .453 | .441 | .736 | .353 | .458 | .495 | .493 | |
e | listwise.likelihood | .749 | .352 | .307 | .416 | .725 | .316 | .479 | .518 | .483 | |
f | pairwise.heapsort | .738 | .359 | .324 | .407 | .744 | .328 | .487 | .543 | .491 | |
g | pairwise.bubblesort | .733 | .363 | .423 | .421 | .756 | .355 | .490 | .550 | .511 | |
h | setwise.heapsort | .752 | .346 | .297 | .402 | .726 | .321 | .473 | .513 | .479 | |
i | setwise.bubblesort | .768 | .346 | .388 | .424 | .754 | .343 | .479 | .534 | .505 |
Methods | NDCG@10 | #Inf. | Pro. tokens | Gen. tokens | Lat.(s) | |
large | heapsort.generate | .670 | 125 | 40461 | 627 | 8 |
heapsort.likelihood | .670 | 125 | 40458 | - | 5 | |
bubblesort.generate | .678 | 461 | 147774 | 2302 | 29 | |
bubblesort.likelihood | .678 | 461 | 147752 | - | 19 | |
xl | heapsort.generate | .693 | 130 | 41666 | 647 | 10 |
heapsort.likelihood | .693 | 130 | 41667 | - | 6 | |
bubblesort.generate | .705 | 467 | 149949 | 2335 | 35 | |
bubblesort.likelihood | .705 | 467 | 149949 | - | 20 | |
xxl | heapsort.generate | .706 | 130 | 42077 | 651 | 20 |
heapsort.likelihood | .706 | 130 | 42071 | - | 17 | |
bubblesort.generate | .711 | 468 | 150765 | 2342 | 73 | |
bubblesort.likelihood | .711 | 468 | 150765 | - | 60 |
@@ 5.2. Efficiency Results
Method | NDCG@10 | TREC DL 2019 | TREC DL 2020 |
pairwise.heapsort | 0.6800 | $3.39 | $3.40 |
setwise.bubblesort | 0.6800 | $4.62 | $4.67 |
pairwise.allpair | 0.6783 | $90.59 | $90.60 |
listwise.likelihood | 0.6745 | $2.83 | $2.86 |
setwise.heapsort | 0.6743 | $1.28 | $1.27 |
pairwise.bubblesort | 0.6550 | $12.28 | $11.89 |
pointwise.yes_no | 0.6398 | $0.48 | $0.49 |
listwise.generation | 0.5929 | $3.49 | $3.75 |
pointwise.qlm | 0.5343 | $0.46 | $0.46 |
Regarding computational and runtime efficiency, the results presented in Table 2 indicate that both Pointwise methods exhibit fewest inference, prompt tokens, and no generated tokens. Furthermore, their computational efficiency and query latency are optimized due to efficient GPU-based batched inference. It is worth noting, however, that these methods do come with certain limitations. Specifically, they require access to the model output logits (thus currently limiting their use to just open source LLMs) and are less effective when used with larger models. In contrast, pairwise.allpair appears to be the most expensive method that consumes the most number of prompt tokens and generated tokens due to the large number of document pair preferences needed to be inferred. Hence, even with GPU batching, pairwise.allpair still has the worst query latency. In constrast, approaches utilizing our Setwise prompting—namely, listwise.likelihood, setwise.heapsort, and setwise.bubblesort, are far more efficient than their counterparts, listwise.generate, pairwise.heapsort, and pairwise.bubblesort respectively. Notably, these improvements are achieved without compromising effectiveness. Section 5.4 will discuss further approaches on improving efficiency.
Table 5 shows calculations for the estimated cost of API calls; this estimation is obtained using the OpenAI GPT-4 cost structure, and applying this same structure to the number of tokens measured in our experiments. At time of writing, OpenAI costs were /1,000 prompt tokens and /1,000 generated tokens. To estimate the token count if GPT-4 were used, we average the number of prompt tokens and generated tokens from Table 2 across Flan-T5 models. The setwise.bottlesort and pairwise.heapsort methods show comparable NDCG@10, but pairwise.heapsort is cheaper. On the other hand, our setwise.heapsort provides a reduction of in cost by only marginally reducing NDCG@10 (a 0.8% loss).
5.3. Impact of using Outputs Logits on Setwise
Similar to Pairwise methods, if the model output logits are accessible, our Setwise approaches can also utilize these logits to estimate the likelihood of the most relevant document label. This approach eliminates the need for token generation, requiring only a single LLM forward inference to yield the output results, thus offering a more efficient process. To assess the impact of incorporating model output logits in our Setwise approaches, we conducted experiments on the TREC DL 2019 dataset, with results presented in Table 4. The findings indicate that using model logits resulted in no change in ranking effectiveness, but did lead to lower query latency. This improvement stems from the absence of generated tokens for likelihood estimation. Hence, we conclude that if access to the model output is available, employing likelihood can further enhance the efficiency for our Setwise approach.
@@ 5.4. Effectiveness and Efficiency Trade-offs
To investigate the trade-off between effectiveness and efficiency inherent in our Setwise approach, we set while truncating the documents in the prompt to tokens222This reduction in document length is necessary to ensure prompt size is not exceeded., respectively. The NDCG@10, along with query latency for all models while varying , is visualized in Figure 3a for the TREC DL datasets. As expected, larger reduces query latency but often degrades effectiveness. Notably, the heap sort algorithm consistently proves more efficient than bubble sort. For instance, with Flan-t5-xl and , heap sort achieves strong NDCG@10 with a query latency of 3 seconds. When compared to the other methods outlined in Table 2, this represents the lowest query latency, except for the Pointwise approaches with Flan-t5-large, albeit with superior ranking effectiveness. It’s worth noting that the ranking effectiveness decline with larger values could also be attributed to the increased truncation of passages. LLMs with extended input length capacity might potentially yield improved ranking effectiveness for larger . This area warrants further exploration in future studies.
@@ 6. Conclusion
Future work should focus on evaluating the Setwise prompting approach on a wider array of LLMs, including LLaMA models (Touvron et al., 2023a, b) as well as the OpenAI LLM APIs. Additionally, recent advanced self-supervised prompt learning techniques (Yang et al., 2023; Fernando et al., 2023) could be used to refine the Setwise approach. We make our code and results publicly available at https://github.com/ielab/llm-rankers.

### v2

@@ 5.1. Effectiveness Results
In regards to ranking effectiveness, it is notable that all LLM-based zero-shot ranking approaches demonstrate a significant improvement over the initial BM25 ranking. The only exception to this trend is the pointwise.qlm approach on DL2019 across all models and DL2020 with the Flan-t5-xxl model. Interestingly, as the LLM size increases, the effectiveness of pointwise.qlm decreases. This finding is particularly unexpected, given the common assumption that larger LLMs tend to be more effective.
On the other hand, pointwise.yes_no method achieved a decent NDCG@10 score with Flan-t5-large when compared to other methods. However, effectiveness also did not increase as model size increased. These unexpected results for both Pointwise methods might be attributed to the requirement of a more refined model output calibration process, ensuring their suitability for comparison and sorting across different documents (Qin et al., 2023).
@@
The Listwise approaches (listwise.generation) are far less effective when tested with Flan-t5-large and Flan-t5-xl. However, listwise.generation shows some improvement with Flan-t5-xxl. These results may be attributed to the fact that generating a ranking list requires fine-grained relevance preferences across multiple documents, a task that may exceed the capabilities of smaller models. In contrast, the listwise.likelihood approach, empowered by our Setwise prompt, markedly enhances the ranking effectiveness of the Listwise approach, even when utilizing smaller models. We acknowledge however that listwise.likelihood requires access to the model output logits, whereas listwise.generation does not. In the case of Pairwise and Setwise approaches, they consistently exhibit good ranking effectiveness across various model sizes and datasets.
In Table 3, we present the zero-shot ranking effectiveness of all methods (with the exception of pairwise.allpair due to its computationally intensive nature) across 8 widely-used BEIR datasets. Notably, we identify several different trends that deviate from observations made on the TREC DL datasets.
@@
Firstly, pointwise.qlm exhibits a slightly higher average NDCG@10 score compared to pointwise.yes_no. Moreover, the effectiveness of pointwise.qlm remains stable even as the model size increases. Secondly, listwise.generation demonstrates comparable effectiveness to listwise.likelihood, with the majority of gains obtained in the Touche dataset, where other methods perform worse. Lastly, both Pairwise and Setwise methods that leverage the bubble sort algorithm consistently demonstrate higher average NDCG@10 compared to when they utilize the heap sort algorithm, regardless of the model size. Overall, the variety of results we observe across different experimental settings shows the importance of not drawing conclusions about effectiveness from single datasets or model sizes.
We note that if the LLM output logits are accessible, our Setwise approaches can also utilize these logits to estimate the likelihood of the most relevant document label. This approach eliminates the need for token generation, requiring only a single LLM forward inference to yield the output results, thus avoiding the generation of unexpected tokens. Surprisingly, in our experiments we find that using model logits for our Setwise approaches resulted in no change in ranking effectiveness when compare to generation, suggesting that the inference of our Setwise approaches that fully relies on token generation is very robust.
@@ 5.2. Efficiency Results
Regarding computational and runtime efficiency, the results presented in Table 2 indicate that both Pointwise methods exhibit fewest inference, prompt tokens, and no generated tokens. Furthermore, their computational efficiency and query latency are optimized due to efficient GPU-based batched inference. It is worth noting, however, that these methods do come with certain limitations. Specifically, they require access to the model output logits (thus currently limiting their use to just open source LLMs) and are less effective when used with larger models. In contrast, pairwise.allpair appears to be the most expensive method that consumes the most number of prompt tokens and generated tokens due to the large number of document pair preferences needed to be inferred. Hence, even with GPU batching, pairwise.allpair still has the worst query latency. In contrast, approaches utilizing our Setwise prompting—namely, listwise.likelihood, setwise.heapsort, and setwise.bubblesort, are far more efficient than their counterparts, listwise.generate, pairwise.heapsort, and pairwise.bubblesort respectively. Notably, these improvements are achieved without compromising effectiveness. Section 5.3 will discuss further approaches on improving efficiency.
The setwise.bottlesort and pairwise.heapsort methods show comparable NDCG@10, but pairwise.heapsort is cheaper. On the other hand, our setwise.heapsort provides a reduction of in cost by only marginally reducing NDCG@10 (a 0.8% loss).
@@ 5.3. Effectiveness and Efficiency Trade-offs
To investigate the trade-off between effectiveness and efficiency inherent in our Setwise approach, we set while truncating the documents in the prompt to tokens, respectively. This reduction in document length is necessary to ensure prompt size is not exceeded. The NDCG@10, along with query latency for all models while varying , is visualized in Figure 3a for the TREC DL datasets. As expected, larger reduces query latency but often degrades effectiveness. Notably, the heap sort algorithm consistently proves more efficient than bubble sort. For instance, with Flan-t5-xl and , heap sort achieves strong NDCG@10 with a query latency of 3 seconds. When compared to the other methods outlined in Table 2, this represents the lowest query latency, except for the Pointwise approaches with Flan-t5-large, albeit with superior ranking effectiveness. It’s worth noting that the ranking effectiveness decline with larger values could also be attributed to the increased truncation of passages. LLMs with extended input length capacity might potentially yield improved ranking effectiveness for larger . This area warrants further exploration in future studies.
@@ 5.4. Sensitivity to the Initial Ranking
5.5. Effectiveness and Costs of other LLMs
In the previous sections, we only used Flan-T5 as the backbone LLM. Flan-T5 is a transformer encoder-decoder model. In this section, to better understand the impact of different models, we investigate popular transformer decoder-only LLMs (open-sourced: llama2-chat-7b222https://huggingface.co/meta-llama/Llama-2-7b-chat-hf (Touvron et al., 2023c), vicuna-13b-v1.5333https://huggingface.co/lmsys/vicuna-13b-v1.5 (Zheng et al., 2023); closed-source: OpenAI gpt-3.5-turbo-1106444https://platform.openai.com/docs/models/gpt-3-5) on DL2019 and DL2020. For open-source models we measure cost in terms of query latency in seconds (s); for closed models we measure the API call costs in US dollars ($).555We exclude Pointwise approaches as they are less effective, and pairwise.bubblesort for gpt-3.5-turbo as it is too costly. Results are presented in the Table 4.
Flan-T5 in our previous results is a better backbone than Llama2 and Vicuna, regardless of ranking method. Notably, Setwise exhibits the best overall performance when considering these models, showcasing its robustness. For gpt-3.5-turbo (closed model), we compared Setwise and Listwise using 10 documents at the time (), as this LLM has a longer input context limit; for Listwise, we set window size 10, step size of 5 and repeat sorting twice for fair comparison with Setwise. Listwise achieves the highest effectiveness; however, Setwise achieves similar effectiveness (no significant differences) but at only half the cost.
TREC DL 2019 | TREC DL 2020 | |||||
# | Methods | NDCG@10 | Cost(s or $) | NDCG@10 | Cost(s or $ ) | |
llama2-chat-7b |
listwise.generation | .508 | 144.9 s | .475c | 143.9 s | |
pairwise.bubblesort | .538ac | 64.3 s | .503ac | 61.5 s | ||
pairwise.heapsort | .465 | 18.3 s | .416 | 17.9 s | ||
setwise.bubblesort | .578abc | 38.2 s | .530ac | 38.5 s | ||
setwise.heapsort | .568c | 10.9 s | .514c | 10.9 s | ||
vicuna-13b |
listwise.generation | .639 | 225.9 s | .608 | 226.3 s | |
pairwise.bubblesort | .612 | 166.6 s | .592 | 169.0 s | ||
pairwise.heapsort | .617 | 46.0 s | .582 | 45.6 s | ||
setwise.bubblesort | .622 | 57.6 s | .602 | 58.8 s | ||
setwise.heapsort | .659bcd | 16.7 s | .583 | 16.4 s | ||
gpt-3.5 |
listwise.generation | .712 | 0.045 $ | 67.2 | 0.046 $ | |
pairwise.heapsort | .694 | 0.171 $ | 65.1 | 0.169 $ | ||
setwise.bubblesort | .699 | 0.084 $ | 65.9 | 0.088 $ | ||
setwise.heapsort | .693 | 0.029 $ | 65.6 | 0.028 $ |
We note that there could be potential data contamination with instruction-tuned LLMs: during the instruction fine-tuning tasks, these models could have been fine-tuned on the MS MARCO or BEIR datasets. Although the document ranking task and the ranking prompts used in this paper are unlikely to be part of the instruction fine-tuning dataset used for these LLMs, we acknowledge that data contamination could still artificially impact the effectiveness of the considered LLMs in ranking tasks. However, since we conducted experiments with different methods based on the same instruction-tuned model, we believe our empirical comparison still offers insights for the practical use of LLM-based re-rankers.
@@ 6. Conclusion
Future work should focus on evaluating the Setwise prompting approach on a wider array of LLMs, including LLaMA models (Touvron et al., 2023a, b) as well as the OpenAI LLM APIs. Additionally, recent advanced self-supervised prompt learning techniques (Yang et al., 2023; Fernando et al., 2023) could be used to refine the Setwise approach.